OpenStudy (kingogcobraxx):

im sorry im bothering you so much after this last week of summer school you wont hear from me again

OpenStudy (kingogcobraxx):

@Jadeishere

OpenStudy (jadeishere):

No, it's fine :)

OpenStudy (kingogcobraxx):

it just seems like im waisting peoples time lately and everyone would be happier if i was gone anyway heres the question

OpenStudy (jadeishere):

No one would be happier, trust me. You're not wasting my time, you're helping me use it wisely. Factor it out >.< What is something that adds to -6 and multiplies to +9 -3. (x-3)^2

OpenStudy (kingogcobraxx):

so a?

OpenStudy (kingogcobraxx):

@Jadeishere

OpenStudy (jadeishere):

Yesh

OpenStudy (jadeishere):

@mayankdevnani Super sorry but I'm helping a few other people with really in-depth topics right now, could you try to help this user Mayank? :/ I'll come back when I can

OpenStudy (mayankdevnani):

what do you think dude?

OpenStudy (kingogcobraxx):

no idea how do you do it?

OpenStudy (mayankdevnani):

\[\large \bf (a+b)^2=a^2+b^2+2ab\] \[\large \bf (a-b)^2=a^2+b^2-2ab\]

OpenStudy (mayankdevnani):

remember ^^^

OpenStudy (kingogcobraxx):

oh gosh

OpenStudy (mayankdevnani):

it is simple not much hard to remember

OpenStudy (mayankdevnani):

can you compare your equation with any one of these two equations?

OpenStudy (kingogcobraxx):

b or d?

OpenStudy (mayankdevnani):

well you're close but not d

OpenStudy (kingogcobraxx):

so b

OpenStudy (mayankdevnani):

well there is a positive sign between them ,and we expand them we didn't get any negative sign in expansion

OpenStudy (mayankdevnani):

right?

OpenStudy (kingogcobraxx):

right

OpenStudy (mayankdevnani):

good so we can eliminate option B and D

OpenStudy (mayankdevnani):

right?

OpenStudy (kingogcobraxx):

right

OpenStudy (kingogcobraxx):

c?

OpenStudy (mayankdevnani):

nuu

OpenStudy (mayankdevnani):

LOL

OpenStudy (kingogcobraxx):

lol well theres my answer

OpenStudy (mayankdevnani):

lmaoo

OpenStudy (mayankdevnani):

it's cheating XD

OpenStudy (kingogcobraxx):

the only one i thought it wasnt XD

OpenStudy (mayankdevnani):

XD

OpenStudy (mayankdevnani):

\[\large \bf (x-3)^2=x^2+(3)^2-2(x)(3)\]

OpenStudy (mayankdevnani):

this would lead us to option A

OpenStudy (mayankdevnani):

what you think?

OpenStudy (kingogcobraxx):

a

OpenStudy (mayankdevnani):

oml

OpenStudy (mayankdevnani):

you're correct mate

OpenStudy (mayankdevnani):

good job

OpenStudy (kingogcobraxx):

@mayankdevnani

OpenStudy (mayankdevnani):

your ideas ?

OpenStudy (kingogcobraxx):

c lol

OpenStudy (mayankdevnani):

*standing ovation*

OpenStudy (mayankdevnani):

correct dude

OpenStudy (kingogcobraxx):

@mayankdevnani i didnt press post ... smh

OpenStudy (kingogcobraxx):

last queestion cmon bro @mayankdevnani

OpenStudy (mayankdevnani):

your ideas ?

OpenStudy (kingogcobraxx):

c

OpenStudy (mayankdevnani):

correct

OpenStudy (kingogcobraxx):

@TheSmartOne

OpenStudy (kingogcobraxx):

@silverspeedfighter

TheSmartOne (thesmartone):

One post per question :P

OpenStudy (kingogcobraxx):

@A_Burning_Masquerade

TheSmartOne (thesmartone):

prime factorization of 120 keep dividing it by primes start with 2, until you can't divide it by 2 anymore then see if it's divisible by 3, then 5, and the next primes until all you're left with are prime numbers